OpenStudy (anonymous):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 11 3x + 4y = 24

OpenStudy (anonymous):

do you have florida virtual school?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh that sucks but so do i, thats how i recognized it because im taking algebra too.

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

do u know what 2 do ?

OpenStudy (anonymous):

yea what module is this? soo maybe i can go to it and show you my answer....but if i do please do write exactly what i did..

OpenStudy (anonymous):

ok it's module 6, 6.10

OpenStudy (anonymous):

2x + y = 11 3x + 4y = 24 take equation 1 and substitute into 2: y = 9 - 2x 3x + 4(9 - 2x) = 26 3x + 36 - 8x = 26 -5x = -10 x = 2 y = 9 - 4 y = 5 (2,5) *i did part 1 not 2 on the test Sorry!, but you can just write whatever

OpenStudy (anonymous):

*i mean i did part 2 not part 1..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea your welcome and if you need any help jusst tag me like this @Snapbacklive because i go on here everyday , because i also need help on flvs,im on module 8(:

OpenStudy (anonymous):

what time do you come on here

OpenStudy (anonymous):

i come like at 11am, and i usually stay till late but idk about tomorrow because i might go out with friends soo yea... but ill try! because i start high school next week and i want to get this over with soo that's why im always on haha(:

OpenStudy (anonymous):

but if you want start on the module 7 pre test now so when you get to 7.01 you can ask me w.e you need help on because im going to stay up late today

OpenStudy (anonymous):

well i have one more question

OpenStudy (anonymous):

sure what is it?

OpenStudy (anonymous):

You graph a system of equations and the two lines intersect one another. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically

OpenStudy (anonymous):

from what module or test is this?

OpenStudy (anonymous):

same one

OpenStudy (anonymous):

ok so if the lines are on top of each each other how many solutions is it?

OpenStudy (radar):

There is an error in your solution @Rkb<3NH

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

which one

OpenStudy (radar):

When he isolated y in the first equation he typoed and used 9 instead of 11

OpenStudy (anonymous):

what do you mean?

OpenStudy (radar):

2x + y = 11 y= 11 - 2x Not y = 9 - 2x

OpenStudy (anonymous):

ohh!! okay im sorry..wait il fix it noww

OpenStudy (radar):

The step by step was a good think as it makes it easy to see an error.

OpenStudy (anonymous):

2x + y = 11 3x + 4y = 24 take equation 1 and substitute into 2: y = 9 - 11x 3x + 4(9 - 11x) = 26 3x + 36 - 44x = 26 -41x = -10 @radar can you help me i got stuck(:

OpenStudy (radar):

2x+y=11 then y = 11 - 2x substituting in the second equation you get 3x + 4(11-2x)=24 3x + 44 -8x = 24 -5x=-20 x=4 did you follow?

OpenStudy (anonymous):

ohh haha yea wow how easy is it too make a simple mistake but thank you(:

OpenStudy (radar):

No problem, it is obvious you know what you're doing, but a typo can really mess u up. Good luck with these.

OpenStudy (anonymous):

Thanks for helping out(:

OpenStudy (anonymous):

how about my question

OpenStudy (radar):

That was part 2 of your question....I think.

OpenStudy (anonymous):

You graph a system of equations and ***the two lines intersect one another.*** Explain what this tells you about the: 1.solution to the system 2.the two equations 3.the result if the system was solved algebraically * for #1: the system will have infinite solutions because the one solution because the lines are perpendicular so they only intersect once.

OpenStudy (radar):

You substitute the value of x (which is 4) into the 2x + y =11 getting 8 + y =11 y=11-8=3 x=4, y=3

OpenStudy (radar):

If the lines do intersect there is one solution for the system. If the lines do not intersect (meaning they are parallel) there is no single solution for that system.

OpenStudy (anonymous):

for #2: the equation could be 2y = 2x + 8 2x + 2y = 4

OpenStudy (anonymous):

for # 3: If the system was solved algebraically, you'll probably end up with, such as 0 = 0.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome(: anymore help?

OpenStudy (anonymous):

no i will have more questions tomorrow, thanks for ur help

OpenStudy (anonymous):

okay (: